OpenStudy (anonymous):

A car gets 24 miles per gallon of gasoline. How many gallons of gasoline would the car need to travel 144 miles?

OpenStudy (mathteacher1729):

So 1 gallon = 24 miles, how many gallons to 48 miles? If you can figure that out, how many gallons to 72 miles? (then just keep going)

OpenStudy (anonymous):

Ok thank you!

OpenStudy (anonymous):

6 gallons of gasoline you divide 144 by 24 then to check your answer multiply your answer by 24

OpenStudy (anonymous):

Think about it this way.\[\text{gallons}=\text{miles}\cdot \frac{\text{gallons}}{\text{miles}}=\boxed{\dfrac{\text{miles}}{\left(\dfrac{\text{miles}}{\text{gallons}}\right)}}\]So just plug in how many miles per gallon the vehicle requires and then how many miles you've gone.

OpenStudy (anonymous):

Couldn't you divide 144 by 24 and get 6 gallons?

OpenStudy (anonymous):

Yep that's exactly the equation I wrote above.